OpenStudy (anonymous):

Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

10A + 6(5) = 80 10A + 30 = 80 (80 - 30 = 50, so we know that $50 more is needed to complete the sentence) 10(5) + 30 = 80 50 + 30 = 80

OpenStudy (anonymous):

So there is 5 adult tickets and 5 children tickets?

OpenStudy (anonymous):

@itzChristopher Thank you for your help. c: